# The Three Regimes of Offline-to-Online Reinforcement Learning

### v2

@@ 5.1 Empirical Values of Stability and Plasticity
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 24 2 1
@@
To ground the concepts in knowledge decomposition (Eq. 6) in practice, we report the empirical values of stability, plasticity, and the improvement for different fine-tuning methods in Table 1 (Superior and Inferior regimes), Table 7 (Comparable regime), and Table 8 (all regimes).
  • Improvement: The highest average improvement is regime-dependent. In the Superior regime, it is achieved by “warmup”, which is a $\pi_{0}$-centric method; in the Inferior regime, it is achieved by “offline data + reset”, which is a $\mathcal{D}$-centric method. In the Comparable regime, the best-performing method is “offline RL + offline data”, with a small advantage. These patterns are consistent with intuition: in the Superior regime, $\pi_{0}$ provides more valuable knowledge to leverage, whereas in the Inferior regime, the offline dataset $\mathcal{D}$ contains the more valuable knowledge.
@@ 5.2 Superior Regime: $J(\pi_{0})>J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ achieves substantially higher performance than the behavior policy, i.e., $J(\pi_{0})>J(\pi_{\mathcal{D}})$. In such cases, the offline dataset offers limited additional value, and the primary concern becomes preserving stability relative to $\pi_{0}$ during online fine-tuning.
@@
Beyond aggregate comparisons, the analysis of specific design choices highlights key trade-offs between stability and plasticity. Within the two variants of $\pi_{0}$-centric methods, online data warmup achieves better performance in 27 out of 32 settings, reflecting its ability to preserve the pretrained policy’s knowledge while maintaining sufficient plasticity for fine-tuning. In contrast, offline RL regularization provides stronger stability, leading to substantially less performance degradation during the early fine-tuning phase. However, this strong stability comes at the cost of reduced plasticity, which limits long-term improvements during fine-tuning, making it better than online data warmup in only 5 out of 32 settings, where the pretrained policy is already close to the optimal policy. The same applies to the combination of offline RL regularization with offline data replay, which exhibits the strongest stability among all methods considered. These contrasts highlight the importance of considering each setting and identifying the the method that best balances the underlying stability–plasticity trade-off.
@@ 5.3 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ performs much worse than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})<J(\pi_{\mathcal{D}})$. Thus, $\pi_{0}$ contributes substantially less useful knowledge than the offline dataset, making it crucial to retain and leverage the offline data during online fine-tuning.
@@
Takeaway: In the Inferior regime, where the pretrained policy $\pi_{0}$ performs substantially worse than the offline dataset $\mathcal{D}$, $\mathcal{D}$-centric methods typically provide more effective fine-tuning than $\pi_{0}$-centric methods. Parameter reset can also be beneficial, particularly when the pretrained agent performs very poorly, as it restores plasticity and allows the agent to adapt more effectively.
@@ 5.4 Comparable Regime: $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy and the behavior policy achieve similar performance, i.e., $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$. Representative results are shown in Figure 5, with the complete results provided in the appendix (Figure 9).
@@
Figure 5: Representative fine-tuning results for the Comparable regime. The first two subplots illustrate cases consistent with our framework’s predictions, while the latter two show mismatches with only small mean differences.
@@
5.5 Analysis of Q-Function During Fine-Tuning
Lastly, because critic learning is central to value-based methods, we examine how the Q-function evolves during fine-tuning in the Superior and Inferior regimes. Figure 6 illustrates a representative pattern using CalQL pretraining, with one Superior regime (kitchen-partial-v0) and one Inferior regime (kitchen-complete-v0). In each regime, we compare a $\pi_{0}$-centric method (warmup) with a $\mathcal{D}$-centric method (offline data).
@@ 6 Conclusion
Our regime taxonomy provides an efficient lens for understanding offline-to-online RL by compressing complex phenomena into a small number of discrete categories. Such taxonomies have been highly successful in both the natural sciences and machine learning, for example in imitation learning, where regimes have been proposed based on density ratios (spencer2021feedback) or dataset size (belkhale2023data). At the same time, we recognize that discretizing behavior into three regimes is a simplification, particularly in situations where return does not adequately capture the usefulness of pretrained policies or offline datasets, for example in long-horizon sparse-reward settings. Real systems often lie along a continuum, and other dataset characteristics, such as coverage, play an important role but remain difficult to capture consistently. Extending the framework to incorporate such dimensions is an important direction for future work.
Our work also connects offline-to-online RL to a broader body of research on the stability–plasticity dilemma in deep learning and neuroscience. Recent studies in continual and online deep RL have focused on characterizing forgetting and plasticity, but applications to the offline-to-online transition have been limited. We show that stability–plasticity is not only a useful lens for analyzing this setting, but also a source of practical guidance: it predicts which design choices are effective in which regimes, moving beyond coining terminology to actionable prescriptions.
@@ Appendix A Related Work
Offline-to-online RL seeks to combine the strengths of offline pretraining with the adaptability of online fine-tuning. Early approaches focused on extending offline RL regularization methods into the online regime, constraining fine-tuning updates to remain close to the pretrained policy. For example, Advantage Weighted Actor-Critic (AWAC) (nair2020awac) and Implicit Q-Learning (IQL) (kostrikov2021offline) applied offline regularization techniques directly to online fine-tuning. Building on the similar idea, PROTO (li2023proto) introduced KL regularization to explicitly constrain the online policy to the pretrained one. Another line of work proposes new replay strategies for incorporating offline data more effectively. lee2022offline propose balanced replay to mitigate distribution shift and bootstrap error when transitioning from offline to online learning, liu2024energy employs a diffusion model to select or generate samples. Alternative strategies separate the roles of exploration and exploitation during fine-tuning. Jump-Start RL (JSRL) (uchendu2023jump), maintains a fixed guided policy from pretraining alongside an exploration policy that is updated online, progressively transferring control from the pretrained policy to the learned one. More recent directions include expanding the action space via policy set expansion (PEX (PEX)) and Bayesian methods for uncertainty-aware exploration (BOORL (hu2024bayesian)). Some approaches skip offline pretraining but still make use of offline data. For example, song2022hybrid and ball2023efficient start training directly with online RL while incorporating offline datasets, providing another way to combine offline data with online interaction.
Plasticity and Stability.
@@
Plasticity and stability have long been recognized as central, often competing, objectives in learning systems. In neuroscience, this tension is formalized as the stability–plasticity dilemma (mermillod2013stability), highlighting the challenge of integrating new knowledge without overwriting previously acquired competencies. In machine learning, similar dynamics manifest when agents must adapt to new data while preserving useful prior knowledge. Early work on continual and lifelong learning addressed this challenge via regularization techniques (kirkpatrick2017overcoming), replay buffers (rolnick2019experience) and modular architectures (rusu2016progressive). More recently, researchers have observed that insufficient plasticity can also hinder online deep RL, motivating methods designed to enhance plasticity of the neural network during training (nikishin2022primacy; sokar2023dormant; dohare2024loss). In our work, we extend this perspective by framing offline-to-online RL explicitly as a stability–plasticity problem, where the central challenge is to balance the preservation of prior knowledge from pretraining with the adaptability needed to learn from new online experiences.
@@ B.1 Offline Pretraining
For offline pretraining, we train CalQL for 1M gradient steps on AntMaze, 20k on Adroit, and 250k on both Kitchen and MuJoCo locomotion tasks. ReBRAC is trained for 1M gradient steps on AntMaze, 100k on Adroit, 250k on Kitchen, and 500k on MuJoCo tasks. For the behavior cloning (BC) baseline, we perform 500K gradient steps of policy learning followed by 100k steps of fitted Q evaluation (voloshin2019empirical) (FQE) to obtain a Q-function for subsequent RL fine-tuning. Since this work primarily focuses on the online fine-tuning stage of offline-to-online RL, we do not modify the offline pretraining algorithm or its default hyperparameters.
@@
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 26 0 1
@@
Fine-tune
$\pi_{0}$-centric >> $\mathcal{D}$-centric 18 9 0
@@ B.3 Hyperparameters
Table 11: ReBRAC hyperparameters used in our experiments. All hyperparameters follow the best hyperparameters reported in tarasov2023revisiting, except for the Kitchen domain, which was not included; for Kitchen, we performed a hyperparameter search and selected the best configuration.
@@ Appendix C Alternative Taxonomies of Three Regimes
While our framework is defined using the returns of the pretrained agents and the dataset, it is also useful to consider alternative taxonomies based on different metrics. Below we present two such taxonomies, derived from Q-functions and behavior-cloning performance, respectively.
C.1 Q-Function-Based Taxonomy

### v3

@@ 5.1 Empirical Values of Stability and Plasticity
+ offline RL + offline data ($\pi_{0}+\mathcal{D}$) $\mathbf{-0.614\pm 0.354}$ $0.440\pm 0.297$ $0.174\pm 0.441$
To ground the concepts in knowledge decomposition (Eq. 3.2) in practice, we report the empirical values of stability, plasticity, and the improvement for different fine-tuning methods in Table 1 (Superior and Inferior regimes), Table 8 (Comparable regime), and Table 9 (all regimes).
  • Improvement: The highest average improvement is regime-dependent. In the Superior regime, it is achieved by “warmup”, which is a $\pi_{0}$-centric method; in the Inferior regime, it is achieved by “offline data + reset”, which is a $\mathcal{D}$-centric method. In the Comparable regime, the best-performing method is “offline RL + offline data”, with a small advantage. These patterns are consistent with intuition: in the Superior regime, $\pi_{0}$ provides more valuable knowledge, whereas in the Inferior regime, the offline dataset $\mathcal{D}$ contains the more valuable knowledge.
@@
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 24 2 1
@@ 5.2 Superior Regime: $J(\pi_{0})>J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ achieves substantially higher performance than the behavior policy, which is common for suboptimal or low-quality datasets. In such cases, the offline dataset offers limited additional value, and the primary concern becomes preserving stability relative to $\pi_{0}$ during online fine-tuning.
@@
Beyond aggregate comparisons, the analysis of specific design choices highlights key trade-offs between stability and plasticity. Within the two variants of $\pi_{0}$-centric methods, online data warmup achieves better performance in 27 out of 32 settings, reflecting its ability to preserve the pretrained policy’s knowledge while maintaining sufficient plasticity for fine-tuning. In contrast, offline RL regularization provides stronger stability, leading to substantially less performance degradation during the early fine-tuning phase. However, this strong stability comes at the cost of reduced plasticity, which limits long-term improvements during fine-tuning, making it better than online data warmup in only 5 out of 32 settings, where the pretrained policy is already close to the optimal policy. The same applies to the combination of offline RL regularization with offline data replay, which exhibits the strongest stability among all methods considered. These contrasts highlight the importance of considering each setting and identifying the method that best balances the underlying stability–plasticity trade-off.
@@ 5.3 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ performs much worse than the behavior policy underlying the offline dataset, which is common for near-expert datasets in sparse-reward domains. Thus, $\pi_{0}$ contributes substantially less useful knowledge than the offline dataset, making it crucial to retain and leverage the offline data during online fine-tuning.
@@
We also provide a deeper investigation into the Q-function in Appendix B.5. The results show that replaying the offline dataset can prevent Q-value divergence in the Inferior regime, effectively anchoring the fine-tuning process, while in the Superior regime, value estimates do not diverge even without the dataset.
Takeaway: In the Inferior regime, where the pretrained policy $\pi_{0}$ performs substantially worse than the offline dataset $\mathcal{D}$, $\mathcal{D}$-centric methods typically provide more effective fine-tuning than $\pi_{0}$-centric methods.
@@ 5.4 Comparable Regime: $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy and the behavior policy achieve similar performance. The complete results are provided in the appendix (Figure 7).
@@
5.5 Alternative Taxonomies of Three Regimes
While our framework defines regimes using raw returns, we additionally evaluate alternative taxonomy criteria to assess robustness. Specifically, we consider regime classification based on: (1) dense reward proxies for sparse-reward domains (Adroit, AntMaze, and Kitchen), (2) learned Q-function values, and (3) Behavioral Cloning (BC) performance. To ensure a fair comparison with raw returns, the dense rewards are constructed without access to privileged task information (e.g., maze layouts) and are used as coarse progress proxies. Full details are provided in Appendix C.
As shown in Table 3, our raw-return-based taxonomy outperforms the alternatives, achieving the best classification accuracy and minimizing opposite mismatches. This suggests that heuristic dense reward proxies, constructed with limited domain knowledge, often diverge from the true objectives of sparse-reward tasks, leading to less reliable regime identification than the sparse raw return that directly defines optimization and evaluation.
Table 3: Results of taxonomies based on alternative metrics beyond raw return, averaged across benchmarks. Dense-reward proxies are applied only to sparse-reward domains.
Domains Metric Accuracy $\uparrow$ Opposite mismatch $\downarrow$
Sparse-reward (27 cases) Raw return (ours) 78% 4%
Dense-reward proxy 65% 11%
All domains (63 cases) Raw return (ours) 71% 5%
Q-function 51% 30%
BC performance 41% 5%
@@ 6 Conclusion
Our regime taxonomy provides an efficient lens for understanding offline-to-online RL by compressing complex phenomena into a small number of discrete categories. Such taxonomies have been widely used in both the natural sciences and machine learning, for example in imitation learning, where regimes have been proposed based on density ratios (Spencer et al., 2021) or dataset size (Belkhale et al., 2023). At the same time, we recognize that discretizing behavior into three regimes is a simplification, particularly in situations where raw return does not adequately capture the usefulness of pretrained policies or offline datasets, for example in long-horizon sparse-reward settings. Real systems often lie along a continuum, and other dataset characteristics, such as coverage, play an important role but remain difficult to capture consistently. Extending the framework to incorporate such dimensions is an important direction for future work.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Offline-to-Online RL
Offline-to-online RL seeks to combine the strengths of offline pretraining with the adaptability of online fine-tuning. Early approaches focused on extending offline RL regularization methods into the online regime, constraining fine-tuning updates to remain close to the pretrained policy. For example, Advantage Weighted Actor-Critic (AWAC) (Nair et al., 2020) and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) applied offline regularization techniques directly to online fine-tuning. Building on the similar idea, PROTO (Li et al., 2023) introduced KL regularization to explicitly constrain the online policy to the pretrained one. Another line of work proposes new replay strategies for incorporating offline data more effectively. Lee et al. (2022) propose balanced replay to mitigate distribution shift and bootstrap error when transitioning from offline to online learning, Liu et al. (2024) employs a diffusion model to select or generate samples. Alternative strategies separate the roles of exploration and exploitation during fine-tuning. Jump-Start RL (JSRL) (Uchendu et al., 2023), maintains a fixed guided policy from pretraining alongside an exploration policy that is updated online, progressively transferring control from the pretrained policy to the learned one. More recent directions include expanding the action space via policy set expansion (PEX (Zhang et al., 2023)) and Bayesian methods for uncertainty-aware exploration (BOORL (Hu et al., 2024)). Some approaches skip offline pretraining but still make use of offline data. For example, Song et al. (2022) and Ball et al. (2023) start training directly with online RL while incorporating offline datasets, providing another way to combine offline data with online interaction.
A.2 Plasticity and Stability
@@
Plasticity and stability have long been recognized as central, often competing, objectives in learning systems. In neuroscience, this tension is formalized as the stability–plasticity dilemma (Mermillod et al., 2013), highlighting the challenge of integrating new knowledge without overwriting previously acquired competencies. In machine learning, similar dynamics manifest when agents must adapt to new data while preserving useful prior knowledge. Early work on continual and lifelong learning addressed this challenge via regularization techniques (Kirkpatrick et al., 2017), replay buffers (Rolnick et al., 2019) and modular architectures (Rusu et al., 2016). More recently, researchers have observed that insufficient plasticity can also hinder online deep RL, motivating methods designed to enhance plasticity of the neural network during training (Nikishin et al., 2022; Sokar et al., 2023; Dohare et al., 2024).
We extend this perspective to offline-to-online RL by framing fine-tuning as a stability–plasticity trade-off between preserving knowledge from pretraining and adapting to new online data. While prior work has examined forgetting and plasticity in continual and online RL, their role in the offline-to-online transition has received limited attention. Our framework shows that stability–plasticity is not only an explanatory lens, but also yields actionable guidance by predicting which design choices are effective in different regimes.
@@ B.1 Offline Pretraining
For offline pretraining, we train CalQL for 1M gradient steps on AntMaze, 20k on Adroit, and 250k on both Kitchen and MuJoCo locomotion tasks. ReBRAC is trained for 1M gradient steps on AntMaze, 100k on Adroit, 250k on Kitchen, and 500k on MuJoCo tasks. For the behavior cloning (BC) baseline, we perform 500K gradient steps of policy learning followed by 100k steps of fitted Q evaluation (Le et al., 2019) (FQE) to obtain a Q-function for subsequent RL fine-tuning. Since this work primarily focuses on the online fine-tuning stage of offline-to-online RL, we do not modify the offline pretraining algorithm or its default hyperparameters.
@@
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 26 0 1
@@
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 18 9 0
@@ B.3 Hyperparameters
Table 12: ReBRAC hyperparameters used in our experiments. All hyperparameters follow the best hyperparameters reported in Tarasov et al. (2023), except for the Kitchen domain, which was not included; for Kitchen, we performed a hyperparameter search and selected the best configuration.
@@ B.4 Compute Details
B.5 Analysis of Q-Function During Fine-Tuning
Since critic learning is central to value-based methods, we examine how the Q-function evolves during fine-tuning in the Superior and Inferior regimes. Figure 8 illustrates a representative pattern using CalQL pretraining, with one Superior regime (kitchen-partial-v0) and one Inferior regime (kitchen-complete-v0). In each regime, we compare a $\pi_{0}$-centric method (warmup) with a $\mathcal{D}$-centric method (offline data).
@@ Appendix C Alternative Taxonomies of Three Regimes
While our framework is defined using the returns of the pretrained agents and the dataset, it is also useful to consider alternative taxonomies based on different metrics. Below we present three such taxonomies, derived from dense reward proxy, Q-functions, and behavior-cloning performance, respectively.
C.1 Dense-Reward-Based Taxonomy
Several domains in our study (AntMaze, Adroit, and Kitchen) feature sparse rewards, where raw episode returns can be an incomplete proxy for the usefulness of prior knowledge. In particular, a pretrained policy may achieve low or zero success rate while still consistently reaching meaningful intermediate states that help exploration and fine-tuning. In such cases, classifying regimes solely based on sparse returns may underestimate the value of pretrained policy.
To address this limitation, we introduce a dense-reward-based taxonomy as a complementary regime classifier. We construct task-specific dense reward proxies using limited environmental knowledge (e.g., distance-to-goal, object proximity, or task progress), without assuming access to oracle dense signals. These dense rewards are used only for regime identification and analysis, not for training.
Dense reward proxy for AntMaze.
We define a simple dense proxy as the negative Euclidean distance between the current agent position $x$ and the goal location $g$:
$r_{\text{maze}}=-\|x-g\|_{2}.$ (6)
While AntMaze contains internal walls that make Euclidean distance an imperfect measure of true task progress, we intentionally avoid using shortest-path signals. This choice mirrors the sparse-reward metric by restricting access to privileged environment information, while still providing a coarse indicator of progress toward the goal.
Dense reward proxy for Adroit.
We construct dense reward proxies for the Pen, Relocate, and Door tasks based on their respective success criteria. In each case, we formulate the reward as a negative squared penalty for deviations from the success thresholds.
  • Pen: Success is defined by an object-goal distance $d<0.075$ and orientation similarity $s>0.95$. We define the proxy as:
    $r_{\text{pen}}=-\max(d-0.075,0)^{2}-\max(0.95-s,0)^{2}.$ (7)
  • Relocate: Success requires the object-target distance $d$ to be less than $0.1$. The proxy is defined as:
    $r_{\text{relocate}}=-\max(d-0.1,0)^{2}.$ (8)
  • Door: The goal is achieved when the door position $p$ exceeds $1.35$. We penalize positions below this threshold:
    $r_{\text{door}}=-\min(p-1.35,0)^{2}.$ (9)
Dense reward proxy for Kitchen.
Kitchen involves a sequence of subtasks, for which we design a dense reward that incentivizes sequential completion. Let $\mathcal{T}$ be the ordered set of subtasks. For each subtask $i\in\mathcal{T}$, the reward component $r_{i}$ is determined by its current status:
[…]
Here $d_{i}$ is the current Euclidean distance to the subtask goal, and $d_{i,\text{init}}$ is the distance recorded when the subtask first became active. The total dense reward is the sum over all subtasks: $r_{\text{kitchen}}=\sum_{i\in\mathcal{T}}r_{i}$.
Accuracy on sparse-reward domains.
Following our primary methodology, we employ a $t$-test to assess whether the performance $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$ differs significantly under the dense reward metric. However, unlike the return-based formulation, we omit the significance margin $\delta$ since the widely varying scales of dense rewards across different domains. Based on this $t$-test, we classify each task into the three regimes. The resulting confusion matrix for the dense-reward taxonomy is presented in Table 15; it correctly identifies the regime in 16 out of 27 cases (59%). For comparison, we evaluate our proposed taxonomy within the same sparse-reward settings using margins of $\delta=0.05$ and $\delta=0$ (Table 16(a) and Table 16(b), respectively). Notably, both configurations achieve an accuracy of 78%, substantially outperforming the dense-reward-based alternative.
Table 15: Confusion matrix of fine-tuning results using dense-reward-based taxonomy in sparse-reward domains. Green cells: correct predictions (16/27); red cells: opposite mismatches (3/27); gray cells: adjacent mismatches (8/27). Overall, dense-reward-based taxonomy achieves 59% correct predictions with 11% opposite mismatches.
dense-reward-based Regime
Superior Comparable Inferior
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 4 0 1
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 2 0 2
$\pi_{0}$-centric << $\mathcal{D}$-centric 2 4 12
Table 16: Confusion matrices of fine-tuning results using raw-return-based taxonomy (ours) in sparse-reward domains. We provide our taxonomy using a margin of $\delta=0.05$ (Left) and $\delta=0$ (Right). Green cells indicate correct predictions, red cells denote opposite mismatches, and gray cells represent adjacent mismatches. In both settings, our taxonomy achieves 78% accuracy (21/27 correct) with only 4% opposite mismatches.
(a) With Margin ($\delta=0.05$)
Pretraining Regime (Ours)
Superior Comparable Inferior
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 4 0 1
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 1 1 2
$\pi_{0}$-centric << $\mathcal{D}$-centric 0 2 16
(b) No Margin ($\delta=0$)
Pretraining Regime (Ours)
Superior Comparable Inferior
Fine-tune $\pi_{0}$-centric >> $\mathcal{D}$-centric 4 0 1
$\pi_{0}$-centric $\approx$ $\mathcal{D}$-centric 2 0 2
$\pi_{0}$-centric << $\mathcal{D}$-centric 0 1 17
C.2 Q-Function-Based Taxonomy